# SCOPE: Stochastic Cartographic Occupancy Prediction Engine for Uncertainty-Aware Dynamic Navigation

### v1

@@ Abstract
This article presents a family of Stochastic Cartographic Occupancy Prediction Engines (SCOPEs) that enable mobile robots to predict the future states of complex dynamic environments. They do this by accounting for the motion of the robot itself, the motion of dynamic objects, and the geometry of static objects in the scene, and they generate a range of possible future states of the environment. These prediction algorithms are software-optimized for real-time performance for navigation in crowded dynamic scenes, achieving 10 times faster inference speed and 3 times less memory usage than the original engines. Three simulated and real-world datasets collected by different robot models are used to demonstrate that these proposed prediction algorithms are able to achieve more accurate and robust stochastic prediction performance than other algorithms. Furthermore, a series of simulation and hardware navigation experiments demonstrate that the proposed predictive uncertainty-aware navigation framework with these stochastic prediction engines is able to improve the safe navigation performance of current state-of-the-art model- and learning-based control policies.
@@ I Introduction
Autonomous mobile robots are beginning to enter people’s lives and try to help us provide different last-mile delivery services, such as moving goods in warehouses or hospitals and assisting grocery shoppers [1, 2, 3]. To realize this vision, mobile robots are required to safely and efficiently navigate through complex and dynamic environments filled not only with static obstacles (e.g., tables, chairs, and walls) but also with many moving people and/or other mobile robots. The first prerequisite for robots to navigate and perform tasks is to use their sensors to perceive the surrounding environment. This work focuses on the next step, which is to accurately and reliably predict how the surrounding environment will change based on these sensor data, as shown in Fig. 1. This will allow to proactively take actions based on their predictions and their prediction uncertainties to avoid potential future collisions, a key part of improving autonomous robot navigation. Note that since this general perception-prediction-control navigation framework is a complex system, it is very important to make these algorithms hardware-friendly, especially for mobile robots with limited resources.
@@
In this article, we propose a family of deep neural network-based Stochastic Cartographic Occupancy Prediction Engines (i.e., SCOPE, SCOPE++, and SO-SCOPE) for resource-constrained mobile robots to provide stochastic future state predictions and enable uncertainty-aware navigation in crowded dynamic scenes. This article is an evolution of our previous conference work [4], in which we defined the architecture of the SCOPE and SCOPE++ predictors, [^1] In our previous work [4] we used the acronym SOGMP (Stochastic Occupancy Grid Map Predictor) instead of SCOPE. validated their ability to predict occupancy grid maps (OGMs), and demonstrated their utility for robotic navigation. The primary extensions of [4] are a more in-depth analysis of SCOPE and SCOPE++, software optimization of the performance of those predictors for real-time usage onboard mobile robots (i.e., SO-SCOPE), and additional experimentation to demonstrate the performance of SO-SCOPE for prediction and navigation. Therefore, integrating our previous work [4] and providing more technical details and more complete experiments, this article presents six primary contributions:
@@
    We design an algorithmic pipeline called SCOPE++ that can use a short history of robot odometry and lidar measurements to predict a distribution of potential future robot/environment states. SCOPE++ includes modules to compensate for ego-motion of the robot, to segment static/dynamic objects in the scene, to predict future scenes using a ConvLSTM network, and to sample other future scenes using a variational autoencoder (VAE).
@@
    We analyze the running time and memory usage of each module of SCOPE++ to identify computational bottlenecks. Based on this, we compress the VAE by performing an in-depth statistical analysis of its statistics and by using knowledge distillation techniques. The resulting software-optimized SCOPE (SO-SCOPE) is able to run in real-time on mobile robot hardware.
@@
    We validate the ability of our SCOPE predictors (i.e., SCOPE++, SCOPE, and SO-SCOPE) to predict OGMs using three OGM datasets, each of which comes from a different robot model. We find that SCOPE has a smaller absolute error, higher structural similarity, higher tracking accuracy, and lower computational resource demands than other state-of-the-art approaches (i.e., ConvLSTM [5], DeepTracking [6], and PhyDNet [7]). We also perform a detailed analysis of the correctness, diversity, and consistency of the uncertainty estimates from SCOPE.
@@
    We propose a costmap-based prediction uncertainty-aware navigation framework to incorporate OGM prediction and its uncertainty information into current existing navigation control policies (model-based or learning-based) to improve their safe navigation performance in crowded dynamic scenes.
@@
    We validate the navigation performance in simulated 3D environments with varying crowd densities and in real-world experiments. We find that the predictive uncertainty-aware navigation framework combined with our proposed SCOPE family can improve the navigation performance and safety of extant control policies relative to state-of-the-art solutions, including a model-based controller [8], a supervised learning-based approach [9], and two deep reinforcement learning (DRL)-based approaches [10, 11].
@@
    We open source the OGM prediction code with the OGM dataset (https://github.com/TempleRAIL/scope) and its predictive uncertainty-aware navigation framework (https://github.com/TempleRAIL/scope_nav).
@@ II-A Environment Prediction
While pure image-based methods are simple and explicitly ignore the dynamic information of the environment, many other DL-based approaches exploit the ego-motion information and motion flow of the environment to improve the OGM prediction accuracy. By using a combination of input placement and recurrent states shifting to compensate for the ego-motion, Schreiber et al. [20] extend their previous image-based works [16, 17] to predict DOGMs in moving ego-vehicle scenarios. By extending the deep tracking framework [6] with a spatial transformer module to compensate for ego-motion, Dequaire et al. [21] propose a gated recurrent unit (GRU)-based deep tracking network with a spatial transformer module (STM) for ego-motion compensation to predict multi-steps future states in stationary and moving vehicle settings. Song et al. [22] propose a GRU-based LiDAR-FlowNet to estimate the forward and backward motion flow between two consecutive OGMs and predict future OGMs. Thomas et al. [23, 24] directly encode spatiotemporal information into the world coordinate frame and propose a 3D-2D feedforward architecture, called DeepSOGM, to predict futures. By considering the ego-motion and motion flow together, Mohajerin et al. [25] first use the standard geometric image transformation to compensate for the ego-motion of the vehicle, then propose a ConvLSTM-based difference learning architecture to extract the motion difference between consecutive OGMs, and finally predict multi-step future states. However, most of these motion-based works are designed for vehicle scenarios, and their network models have a large memory footprint and require significant computation, putting them out of the range for small mobile robots with limited resources. Furthermore, all the above-described works only provide deterministic OGM predictions and cannot estimate the uncertainty of future states, which is a key point in helping robots operate in complex, dynamic, and hazardous environments. To address this issue, we propose two versions of a variational autoencoder (VAE)-based stochastic OGM predictor for resource-constrained mobile robots, namely SCOPE++ and SCOPE, both of which predict a distribution of possible future states of dynamic scenes.
@@ II-B Deep Neural Network Optimization
As summarized in surveys [26, 27], the mainstream software acceleration methods can be divided to three categories: 1) network pruning, 2) weight quantization and sharing, and 3) knowledge distillation. The key idea of network pruning is to analyze the different components of complex deep neural networks and remove unimportant components, such as unimportant channels [28, 29], unimportant kernel filters [30, 31], unimportant network connections [32, 33], and even unimportant layers [34, 35]. However, network pruning requires expertise and experience to analyze the network model and can easily lead to performance degradation. To avoid the extra knowledge of pruning, many researchers focus on network model weight quantization and sharing, such as using Huffman coding to quantize weights [36], reduce the number of bits representing model weights [37, 38], and share weights between different network connections or different layers [39, 40]. However, weight quantization often requires a good trade-off between performance and bit quantization, and requires hardware computing devices to support its low-precision arithmetic operations. To keep network model performance and compress neural network, Hinton et al. [41] first proposed the knowledge distillation technique, the key idea of which is to train a small and simple “student” network model to learn the output of a large and complex “teacher” network model. Following the basic distillation idea, many improvement variants are proposed to reduce the learning gap between the “student” network and the “teacher” network such as adding a middle size network as the “teaching assistant” [42, 43], using early-stopped knowledge distillation “teacher” network [44], and incorporating cross-domain distillation training [45]. Although knowledge distillation requires the selection of appropriate “student” or “teacher” networks, it provides more freedom to choose different compressed network architectures and provides the potential possibility of obtaining compressed “student” network models with better generalization capabilities. Therefore, we chose to combine the knowledge distillation technique with the uncertainty quantification module to software-optimize our proposed variational autoencoder (VAE)-based stochastic OGM predictors and make them hardware-friendly for resource-constrained robots.
@@ II-C Uncertainty-Aware Navigation
As we previously discussed, while most approaches [8, 10, 9, 11] use the past or current environment states (i.e., raw sensor data and preprocessed data representations) as perception information for their navigation control policies, many other methods [46, 47, 48, 49, 50, 51] believe that predicted future states can help robots better avoid collisions with pedestrians, and start utilizing the predicted environment states (e.g., pedestrian trajectory prediction and latent state prediction) to improve navigation performance through pedestrian crowds. However, most prediction-based navigation policies assume the outcomes of the interactive environments and their control actions are deterministic. Due to the stochastic nature of the robot and its surrounding environments, there are many uncertainties in environment prediction and control actions. To leverage the uncertainty information in the environmental future states and provide robust and reliable navigation behavior, Kahn et al. [52] propose a collision prediction network to estimate the future collision uncertainty and incorporate the predicted collision uncertainty into Model Predictive Control (MPC) framework to enable uncertainty-aware navigation behavior. Similarly, under the MPC framework, Lütjens et al. [53] use an ensemble of LSTM networks with dropout and bootstrapping to estimate prediction collision probabilities and provide uncertainty-aware navigation around pedestrians. Tang et al. [54] use uncertainty-aware potential field to process prediction uncertainty and incorporate it into MPC framework to provide predictive uncertainty-aware navigation. Furthermore, Sekiguchi et al. [55] convert the uncertainty information of predicted human trajectories into reliability values and incorporate it into a nonlinear MPC framework to realize safe human-followings. However, all these uncertainty-aware works focus on applying their prediction uncertainty information into the MPC framework and cannot be directly adapted to the currently existing model-based and learning-based control policies.
@@
Figure 2: System architectures of the SCOPE++ predictor, SCOPE predictor, and its software-accelerated SO-SCOPE predictor (note that SCOPE omits the Static Objects block compared to SCOPE++). The basic process of the SCOPE++ predictor is: 1) based on a history of robot states, the robot transfers the lidar measurement history to the predicted coordinate frame of the robot to compensate for the ego-motion, 2) these compensated lidar measurements are used to generate a local environment map to account for static objects, and a set of OGMs to account for dynamic objects, and 3) the local map of static objects and the predicted OGM of dynamic objects are fed into an variational autoencoder to predict the future OGM. To accelerate the SCOPE++ predictor, we first follow the SCOPE network architecture and replace the VAE network with single convolutional layer, then use knowledge distillation technology to train the SO-SCOPE network to obtain the prediction information, and finally we model and quantify the prediction uncertainty of the SCOPE++ to obtain uncertainty statistics and use them to generate uncertainty estimates of SO-SCOPE.
@@ III Stochastic Cartographic Occupancy Prediction Engine
This section begins by formulating the problem of OGM prediction in dynamic scenes with many static and dynamic objects. It will then go on to describe our VAE-based predictor, called the Stochastic Cartographic Occupancy Prediction Engine (SCOPE), describing the system architecture as well as how to handle robot motion, pedestrian motion, and static obstacles in dynamic environments.
@@ III-A Problem Formulation
We consider the problem of an unmanned ground vehicle (UGV) moving through an environment filled with both static infrastructure (e.g., walls) and dynamic objects (e.g., people). As the robot moves, we assume that it is able to obtain accurate estimates of its own relative pose and velocity from its odometry sensors or other localization algorithms over short periods of time (on the order of 1 s). We denote the pose and velocity of the robot at time $t$ by $\mathbf{x}_{t}=[x_{t}\ y_{t}\ \theta_{t}]^{T}$ and $\mathbf{u}_{t}=[v_{t}\ w_{t}]^{T}$ respectively. We assume the robot is equipped with a 2D lidar sensor. Let $\mathbf{y}_{t}=[r_{t}\ b_{t}]^{T}$ denote the lidar measurements (range $r$ and bearing $b$) at time $t$.
We assume the world is 2.5D, as is commonly done in mobile robotics applications. Thus, we can represent the environment using a 2D occupancy grid map (OGM). Let $\mathbf{o}_{t}$ denote the OGM at time $t$. We will consider OGMs that are both binary (i.e., each cell is either occupied or free) and probabilistic (i.e., each cell has a probability of being occupied).
Given a history of $\tau$ lidar measurements $\mathbf{y}_{t-\tau:t}$ and robot states $\{\mathbf{x}_{t-\tau:t},\mathbf{u}_{t-\tau:t}\}$, we consider the problem of predicting the future state (i.e., OGM) of the environment $\mathbf{o}_{t+1}$. For notational compactness, let $\mathbf{d}_{t-\tau:t}=\{\mathbf{x}_{t-\tau:t},\mathbf{u}_{t-\tau:t},\mathbf{y}_{t-\tau:t}\}$ denote the history of data over the given window. We can formulate this as a prediction model:
@@
Note that in this paper, we set the lidar history to $\tau=10$, the sampling rate is 10 Hz, and limit the physical size of OGMs to $[0,6.4]$ m along the x axis (forward) and $[-3.2,3.2]$ m along the y axis (left) in the robot’s local coordinate frame. We use a cell size of 0.1 m, resulting in $64\times 64$ OGMs. These settings are consistent with other works on mobile robot navigation  [59]. All data $\mathbf{o},\mathbf{u},\mathbf{y}$ are represented in the local coordinate frame of the robot.

### v2

@@ Abstract
This article presents a family of Stochastic Cartographic Occupancy Prediction Engines (SCOPEs) that enable mobile robots to predict the future states of complex dynamic environments. They do this by accounting for the motion of the robot itself, the motion of dynamic objects, and the geometry of static objects in the scene, and they generate a range of possible future states of the environment. These prediction engines are software-optimized for real-time performance for navigation in crowded dynamic scenes, achieving up to 89 times faster inference speed and 8 times less memory usage than other state-of-the-art engines. Three simulated and real-world datasets collected by different robot models are used to demonstrate that these proposed prediction algorithms are able to achieve more accurate and robust stochastic prediction performance than other algorithms. Furthermore, a series of simulation and hardware navigation experiments demonstrate that the proposed predictive uncertainty-aware navigation framework with these stochastic prediction engines is able to improve the safe navigation performance of current state-of-the-art model- and learning-based control policies.
@@ I Introduction
Autonomous mobile robots are beginning to enter people’s lives and are trying to help us provide different last mile delivery services, such as moving goods in warehouses or hospitals and assisting grocery shoppers [1, 2, 3]. To realize this vision, mobile robots are required to safely and efficiently navigate through complex and dynamic environments filled not only with static obstacles (e.g., tables, chairs, and walls) but also with many moving people and/or other mobile robots. The first prerequisite for robots to navigate and perform tasks is to use their sensors to perceive the surrounding environment. This work focuses on the next step, which is to accurately and reliably predict how the surrounding environment will change based on these sensor data, as shown in Fig. 1. This will allow a robot to proactively act based on its predictions and the associated uncertainty to avoid potential future collisions, a key part of improving autonomous robot navigation. Note that since this general perception-prediction-control navigation framework is a complex and resource-intensive system, it is very important to make these algorithms hardware-friendly (e.g., using smaller computational power, memory usage, and storage usage) and run in real-time, especially for mobile robots with limited resources. A well-performing predictor is useless for practical robotics applications if it consumes a lot of memory and/or cannot run in real-time on a resource-limited robot.
@@
In this article, we propose a family of deep neural network (DNN)-based Stochastic Cartographic Occupancy Prediction Engines (i.e., SCOPE, SCOPE++, and SO-SCOPE) for resource-constrained mobile robots to provide stochastic future state predictions and enable uncertainty-aware navigation in crowded dynamic scenes. This article is an evolution of our previous work [4], in which we defined the architecture of the SCOPE and SCOPE++, [^1] In our previous work [4] we used the acronym SOGMP (Stochastic Occupancy Grid Map Predictor) instead of SCOPE. validated their ability to predict occupancy grid maps (OGMs), and demonstrated their utility for robotic navigation. The primary extensions to [4] include: 1. providing comprehensive OGM prediction performance and resource usage evaluations that include three new state-of-the-art algorithms, 2. mathematically modeling the output characteristics of the prior work to understand its performance, particularly around the statistical distribution of the VAE, 3. designing a novel software-optimized SO-SCOPE predictor that significantly improves the inference speed and addresses the memory-intensive nature of the previous sampling-based SCOPE and SCOPE++ predictors, 4. extending the previous predictive uncertainty-aware navigation framework to enable it to be integrated with resource-intensive learning-based control policies, and 5. providing additional simulation experiments and real-world experiments to demonstrate the performance of these algorithms.
Therefore, integrating our previous work [4] and providing more technically improved materials and more complete experiments, this article presents six primary contributions:
@@
    We design an algorithmic pipeline called SCOPE++ that can use a short history of robot odometry and lidar measurements to predict a distribution of potential future robot/environment states. SCOPE++ includes modules to compensate for the ego-motion of the robot, to segment static/dynamic objects in the scene, to predict future scenes using a ConvLSTM network, and to sample other future scenes using a variational autoencoder (VAE).
@@
    We analyze the running time and memory usage of each module of SCOPE++ to identify computational bottlenecks. Based on this, we compress the VAE by performing an in-depth statistical analysis of its output and by using knowledge distillation techniques. The resulting software-optimized SCOPE (SO-SCOPE) achieves slightly better performance while consuming less memory, performing faster inference, and running in real-time with other resource-intensive algorithms on resource-constrained mobile robot hardware.
@@
    We validate the ability of our SCOPE predictors (i.e., SCOPE++, SCOPE, and SO-SCOPE) to predict OGMs using three OGM datasets (each of which comes from a different robot model) and provide a comprehensive benchmark of prediction performance and resource usage using six state-of-the-art algorithms. We find that the SCOPE family achieves smaller absolute errors, higher structural similarity, higher tracking accuracy, and lower computational resource requirements than other state-of-the-art methods (i.e., ConvLSTM [5], DeepTracking [6], PhyDNet [7], SAAConvLSTM [8], TAAConvLSTM [8], and LOPR [9]). We also perform a detailed analysis of the correctness, diversity, and consistency of the uncertainty estimates from the SCOPE family.
@@
    We propose a costmap-based predictive uncertainty-aware navigation framework to incorporate OGM prediction and its uncertainty information into current existing navigation control policies to improve their safe navigation performance in crowded dynamic scenes.
@@
    We validate the navigation performance in simulated 3D environments with varying crowd densities and real-world experiments. We find that the predictive uncertainty-aware navigation framework combined with our proposed SCOPE family can improve the navigation performance and safety of extant control policies relative to state-of-the-art solutions, including a model-based controller [10], a supervised learning-based approach [11], and two deep reinforcement learning (DRL)-based approaches [12, 13].
@@
    We open source the OGM prediction code with the OGM dataset [14] (https://github.com/TempleRAIL/scope) and its predictive uncertainty-aware navigation framework (https://github.com/TempleRAIL/scope_nav).
@@ II-A Environment Prediction
While pure image-based methods are simple and explicitly ignore the dynamic information of the environment, many other DL-based approaches exploit the ego-motion information and motion flow of the environment to improve the OGM prediction accuracy. By using a combination of input placement and recurrent states shifting to compensate for the ego-motion, Schreiber et al. [22] extend their previous image-based works [19, 20] to predict DOGMs in moving ego-vehicle scenarios. Dequaire et al. [23] extend the deep tracking framework [6] and propose a gated recurrent unit (GRU)-based deep tracking network with a spatial transformer module (STM) for ego-motion compensation to predict multi-steps future states in stationary and moving vehicle settings. Song et al. [24] propose a GRU-based LiDAR-FlowNet to estimate the forward and backward motion flow between two consecutive OGMs and predict future OGMs. Thomas et al. [25, 26] directly encode spatiotemporal information into the world coordinate frame and propose a 3D-2D feedforward architecture, called DeepSOGM, to predict futures. By considering the ego-motion and motion flow together, Mohajerin et al. [27] first use the standard geometric image transformation to compensate for the ego-motion of the vehicle, then propose a ConvLSTM-based difference learning architecture to extract the motion difference between consecutive OGMs, and finally predict multi-step future states. However, most of these motion-based works are designed for vehicle scenarios, and their network models are memory-intensive and computationally intensive, putting them out of the range of resource-limited mobile robots. Furthermore, all the above-described works only provide deterministic OGM predictions and cannot estimate the uncertainty of future states, which is a key point in helping robots operate in dynamic hazardous environments. To address this issue, we propose VAE-based stochastic OGM predictors for resource-constrained robots, namely SCOPE++ and SCOPE, both of which predict a distribution of possible future states of dynamic scenes.
@@ II-B Deep Neural Network Optimization
As summarized in surveys [28, 29], the main methods of software acceleration can be divided into three categories: 1) network pruning, 2) weight quantization and sharing, and 3) knowledge distillation. The key idea of network pruning is to analyze the different components of complex DNNs and remove unimportant components, such as unimportant channels [30, 31], unimportant kernel filters [32, 33], unimportant network connections [34, 35], and even unimportant layers [36, 37]. However, network pruning requires expertise and experience to analyze the network model and can easily lead to performance degradation. To avoid the extra knowledge of pruning, many researchers focus on network model weight quantization and sharing, such as using Huffman coding to quantize weights [38], reduce the number of bits representing model weights [39, 40], and share weights between different network connections or different layers [41, 42]. However, weight quantization often requires a good trade-off between performance and bit quantization, and requires computing devices to support its low-precision arithmetic operations. To keep network model performance and compress neural network, Hinton et al. [43] first proposed the knowledge distillation technique, the key idea of which is to train a small and simple “student” network to learn the output of a large and complex “teacher” network. Following the basic distillation idea, many improvement variants are proposed to reduce the learning gap between the “student” network and the “teacher” network such as adding a middle-size network as the “teaching assistant” [44, 45], using early-stopped knowledge distillation “teacher” network [46], and incorporating cross-domain distillation training [47]. Although knowledge distillation requires the selection of appropriate “student” or “teacher” networks, it provides more freedom to choose different compressed network architectures and provides the potential possibility of obtaining compressed “student” networks with better generalization capabilities. Therefore, we chose to combine the knowledge distillation technique with the uncertainty quantification module to software-optimize our proposed VAE-based stochastic OGM predictors and make them hardware-friendly for resource-constrained robots.
@@ II-C Uncertainty-Aware Navigation
As we previously discussed, while most approaches [10, 12, 11, 13] use the past or current environment states (i.e., raw sensor data and preprocessed data representations) as perception information for their navigation control policies, many methods [48, 49, 50, 51, 52, 53] believe that predicted future states can help robots better avoid collisions with pedestrians, and start utilizing the predicted environment states (e.g., pedestrian trajectory prediction and latent state prediction) to improve navigation performance through pedestrian crowds. However, most prediction-based navigation policies assume the outcomes of the interactive environments and their control actions are deterministic. Due to the stochastic nature of the robot and its surrounding environments, there are many uncertainties in environment prediction and control actions. To leverage the uncertainty information in the environmental future states and provide robust and reliable navigation behavior, Kahn et al. [54] propose a collision prediction network to estimate the future collision uncertainty and incorporate the predicted collision uncertainty into Model Predictive Control (MPC) framework to enable uncertainty-aware navigation behavior. Similarly, under the MPC framework, Lütjens et al. [55] use an ensemble of LSTM networks with dropout and bootstrapping to estimate prediction collision probabilities and provide uncertainty-aware navigation around pedestrians. Tang et al. [56] use the uncertainty-aware potential field to process prediction uncertainty and incorporate it into the MPC framework to provide predictive uncertainty-aware navigation. Furthermore, Sekiguchi et al. [57] convert the uncertainty information of predicted human trajectories into reliability values and incorporate it into a nonlinear MPC framework to realize safe human-followings. However, all these uncertainty-aware works focus on applying their prediction uncertainty information into the MPC framework and cannot be directly adapted to the currently existing model-based and learning-based control policies.
@@
Figure 2: System architectures of the SCOPE++ predictor, SCOPE predictor, and its software-accelerated SO-SCOPE predictor (note that SCOPE omits the Static Objects block compared to SCOPE++). The basic process of the SCOPE++ predictor is: 1) based on a history of robot states, the robot transfers the lidar measurement history to the predicted coordinate frame of the robot to compensate for the ego-motion, 2) these compensated lidar measurements are used to generate a local environment map to account for static objects, and a set of OGMs to account for dynamic objects, and 3) the local map of static objects and the predicted OGM of dynamic objects are fed into an variational autoencoder to predict the future OGM. To accelerate the SCOPE++ predictor, we first follow the SCOPE network architecture and replace the VAE network with a single convolutional layer, then use knowledge distillation technology to train the SO-SCOPE network to obtain the prediction information, and finally, we model and quantify the prediction uncertainty of the SCOPE++ to obtain uncertainty statistics and use them to generate uncertainty estimates of SO-SCOPE.
@@ III Stochastic Cartographic Occupancy Prediction Engine
This section begins by formulating the problem of OGM prediction in dynamic scenes. It will then go on to describe our VAE-based predictors (SCOPE series), describing the system architecture as well as how to handle robot motion, pedestrian motion, and static obstacles in dynamic environments.
@@ III-A Problem Formulation
We consider the problem of a mobile robot moving through an environment filled with both static objects (e.g., walls) and dynamic objects (e.g., people). As the robot moves, we assume that it can obtain accurate estimates of its relative pose and velocity from odometry sensors or other localization algorithms in a short period of time (on the order of 1 s). We denote the pose and velocity of the robot at time $t$ by $\mathbf{x}_{t}=[x_{t}\ y_{t}\ \theta_{t}]^{T}$ and $\mathbf{u}_{t}=[v_{t}\ w_{t}]^{T}$ respectively. We assume the robot is equipped with a 2D lidar sensor. Let $\mathbf{y}_{t}=[\mathbf{r}_{t}\ \mathbf{b}_{t}]^{T}$ denote the lidar measurements (range $\mathbf{r}$ and bearing $\mathbf{b}$) at time $t$. We assume the world is 2.5D, as is commonly done in mobile robotics applications. Thus, we can represent the environment using a 2D occupancy grid map (OGM). Let $\mathbf{o}_{t}$ denote the OGM at time $t$. We will consider OGMs that are both binary (i.e., each cell is either occupied or free) and probabilistic (i.e., each cell has a probability of being occupied).
Given a history of $\tau$ lidar measurements $\mathbf{y}_{t-\tau:t}$ and robot states $\{\mathbf{x}_{t-\tau:t},\mathbf{u}_{t-\tau:t}\}$, we consider the problem of predicting the distribution of future states (i.e., OGMs) of the environment $\mathbf{o}_{t+1}$. For notational compactness, let $\mathbf{d}_{t-\tau:t}=\{\mathbf{x}_{t-\tau:t},\mathbf{u}_{t-\tau:t},\mathbf{y}_{t-\tau:t}\}$ denote the history of data over the given window. We can formulate this as a prediction model:
@@
Note that we set the lidar history to $\tau=10$, the sampling rate is 10 Hz, and limit the physical size of OGMs to $[0,6.4]$ m along the x-axis (forward) and $[-3.2,3.2]$ m along the y-axis (left) in the robot’s local coordinate frame. We use a cell size of 0.1 m, resulting in $64\times 64$ OGMs. These settings are consistent with other works on mobile robot navigation  [61]. All data $\mathbf{o},\mathbf{u},\mathbf{y}$ are represented in the robot’s local coordinate frame.